# Formatting Instructions for ICLR 2025 Conference Submissions

### v1

@@ Abstract
The high computational costs of large language models (LLMs) have led to a flurry of research on LLM compression, via methods such as quantization, sparsification, or structured pruning. A new frontier in this area is given by dynamic, non-uniform compression methods, which adjust the compression levels (e.g., sparsity) per-block or even per-layer in order to minimize accuracy loss, while guaranteeing a global compression threshold. Yet, current methods rely on heuristics for identifying the “importance” of a given layer towards the loss, based on assumptions such as error monotonicity, i.e. that the end-to-end model compression error is proportional to the sum of layer-wise errors. In this paper, we revisit this area, and propose a new and general approach for dynamic compression that is provably optimal in a given input range. We begin from the motivating observation that, in general, error monotonicity does not hold for LLMs: compressed models with lower sum of per-layer errors can perform worse than models with higher error sums. To address this, we propose a new general evolutionary framework for dynamic LLM compression called EvoPress, which has provable convergence, and low sample and evaluation complexity. We show that these theoretical guarantees lead to highly competitive practical performance for dynamic compression of Llama, Mistral and Phi models. Via EvoPress, we set new state-of-the-art results across all compression approaches: structural pruning (block/layer dropping), unstructured sparsity, as well as quantization with dynamic bitwidths. Our code is available at https://github.com/IST-DASLab/EvoPress.
@@ 1 Introduction
Model compression has become a standard way of reducing the deployment costs of large language models (LLMs). Current post-training compression techniques can be roughly categorized into quantization-based, which reduce the bit-width of weights or activations, e.g. (Frantar et al., 2022; Lin et al., 2023; Dettmers & Zettlemoyer, 2022; Tseng et al., 2024), pruning-based, which sparsify the weight matrices, e.g. (Frantar & Alistarh, 2023; Yin et al., 2024), or structured pruning / layer dropping, which drop entire model components, e.g. (Kim et al., 2024; Men et al., 2024). While constantly improving their performance, existing compression methods are reaching diminishing returns in terms of accuracy-vs-compression (Dettmers et al., 2023; Tseng et al., 2024).
In this context, a new direction is dynamic, or non-uniform, layer-wise compression, in which different layers can be compressed to various levels, according to their “sensitivity” relative to the model output. Dynamic compression allows to maximize model accuracy while satisfying a given compression requirement, e.g. a target model size. Instance-specific solutions for this problem have already been proposed for essentially every compression type: sparsity (Yin et al., 2024), quantization (Frantar & Alistarh, 2022), or layer dropping (Kim et al., 2024; Men et al., 2024). Broadly, these approaches work by assigning an error/sensitivity score to each layer and compression level, which measures the impact of its compression on output loss increase. Then, one calculates a compression assignment which minimizes the sum of error scores, while still satisfying the global compression constraint. Thus, such approaches inherently assume error monotonicity: i.e., that a lower sum of error scores implies a lower compression error for the entire model.
Our work starts from the observation that error monotonicity does not hold generally for LLM compression: specifically, there are instances where compressed models with lower sums of per-layer errors can perform worse than models with higher error. We illustrate this fact in Table 1, which shows an instance of a layer dropping configuration where keeping more blocks leads to massively higher perplexity than an instance which prunes strictly less blocks.
@@
Table 1: Depth pruning is not monotone. In this example (Llama-3-8B with Fineweb-Edu calibration), removing strictly more blocks (depicted in orange) can improve perplexity across sources. Left half of block corresponds to attention layer, right half to MLP.
[…]
Llama-3-8B                                                                                                                           5.54 8.80 7.72
                                                                                                                          188.01 147.25 70.46
                                                                                                                          24.39 35.53 26.24
Our algorithm guarantees convergence: specifically, any linear fitness function defined on the $n$-dimensional hypercube will be maximized in expected $O(k(n-k)/\lambda)$ generations under the constraint $\|x\|_{1}=k$, where $\lambda$ is the number of offspring. The proof is quite non-trivial, as it needs to adapt stochastic drift analysis techniques to the case where multiple offspring are examined in each sub-step. In Figure 1, we illustrate the algorithm’s fast convergence and high efficiency on a practical example with correlated block dropping on Llama-3-8B, where we determined the optimum via (expensive) exhaustive search: EvoPress is able to reach the optimum in only $6$ generations, using a total of only $56$ model evaluations. A key advantage of our approach is that it is agnostic of the model architecture and compression type. We illustrate this via experiments, which are the first to span all three compression methods, across different LLM families. Specifically, results show that EvoPress significantly improves upon all prior work on depth pruning in terms of accuracy-vs-compression, especially at medium levels, and also outperforms the prior best methods – OWL and dynamic programming, respectively – for non-uniform pruning and quantization. Moreover, it can do so efficiently: the full version of EvoPress, applied at high compression granularity, will converge in a few hours on a single RTX 3090 GPU, and we also present a lightweight version which utilizes fewer samples and converges in $\sim$ 1 hour in the same setting, on an 8B-parameter model.
@@ 2 Related Work
To our knowledge, we are the first to present a unified approach which covers all types of post-training LLM compression (i.e., layer dropping / depth pruning and non-uniform pruning / quantization) – so far, these problems have generally been approached independently.
Depth Pruning. Recently, there has been a lot of interest in compression by removing entire transformer blocks, both for efficiency and to gain insights about the language model itself. Most methods are based on scoring the importance of each block, and then maximizing the importance of the resulting model by removing the blocks of lowest importance. Weight Subcloning (Samragh et al., 2023) proposed a multi-step process to find good initializations for an untrained smaller model given an already trained larger one, where the importance of each block is scored based on the ratio of $\ell_{2}$ norms between the output embeddings of the block with and without the residual connection. Shortened Llama (Kim et al., 2024) proposes scoring each block by measuring the perplexity after removing the respective block from the full model. ShortGPT (Men et al., 2024) uses the cosine similarity between the input and output embeddings of each block to assess its importance. By contrast, Gromov et al. (2024) restrict themselves to removing consecutive blocks and score each of these removal configurations using cosine similarity.
@@
The recent OWL method (Yin et al., 2024) focuses on non-uniform pruning of LLMs, and provides consistent improvements over uniform profiles via a layer scoring system which analyzes the activation outlier structure, but does not have any theoretical guarantees. Experimentally, we find that OWL is effective especially for Llama-family models (Touvron et al., 2023) and at moderate sparsities, but observe significant gaps in favor of EvoPress across all models and compression levels.
@@
Random search is also popular in the context of structural pruning and Neural Architecture Search (NAS) (Chen et al., 2020; Dong et al., 2021; Wang et al., 2020; Xu et al., 2021; Yin et al., 2021; Molchanov et al., 2022; Kurtić et al., 2024). However, such methods also rely heavily on re-training and have notoriously high costs, which limits their applicability to post-training compression of LLMs. Due to its low sample complexity, we believe that EvoPress could be extensible to lightweight NAS as well, and plan to investigate this in future work.
@@ 3 Method
where $f(\hat{M})$ quantifies the performance of the compressed model $\hat{M}$ and $g(\hat{M})$ represents the compression constraint. For simplicity, we will define $g$ as the model’s total size (in terms of parameters); however, the proposed method can be readily adapted to accommodate other practical constraints, such as inference speed.
@@
Given the specified database, any compressed model is completely characterized by its compression level for each unit (per layer or per block). With $n$ units, each available in $m$ compression levels, our objective is to find
@@
where we are searching over the set of $n$-tuples over $[m]$. Assessing the performance of a model in practice typically involves benchmark tasks, which have limited scope and require lengthy evaluation. We address these challenges by using the base model as the gold standard and focusing solely on the relative degradation of our compressed models. To quantify this degradation, we measure the Kullback-Leibler (KL) divergence between the two models, as it has proven particularly robust with limited data. Empirically, we observed that already around $65536$ tokens of calibration data (corresponding to $8$ full sample sequences for Llama-3-8B) are sufficient to reliably determine the quality of the lightweight model. To avoid confusion, we will refrain from inverting the fitness function and from now on consider the minimization problem
$\hat{M}^{*}=\operatorname*{arg\,min}_{v\in[m]^{n}}D_{KL}(P_{M}\parallel Q_{\hat{M}_{v}})\quad\text{subject to}\quad g(\hat{M_{v}})\leq C,$
@@
[…]
       // Only for non-integer target compression
[…]
@@
[…]
Algorithm 1 EvoPress: A $(1+\lambda)$-Evolutionary Algorithm with Level-Switch Mutation and Multi-Step Selection for Maximizing a Fitness Function $f:[m]^{n}\rightarrow\mathbb{R}$
@@
For initialization, we apply the target level directly if it matches an available setting (e.g., all layers at 70% sparsity for an average of 70% sparsity). If the target falls between two compression levels (e.g., for block dropping), we initialize by randomly sampling candidates with some units compressed to the next lower level, and others to the next higher level, selecting the fittest among them. A high level overview of this optimization procedure can be found in Algorithm 1.
@@
Crossover, for instance, is only effective if population diversity is preserved, for example measured by the sum of pairwise Hamming distances between individuals (Jansen & Wegener, 2002; Opris et al., 2024). While this promotes a more thorough exploration of the search space, it requires allocating resources to less promising regions, slowing down the progress toward optimal solutions. Similarly, non-elitist algorithms, despite their ability to escape local optima (Dang et al., 2021; Jorritsma et al., 2023; Lengler et al., 2024), also incur costs by frequently discarding potentially useful individuals. Consequently, these approaches should be reserved for situations where the fitness landscape is highly rugged, and escaping local optima is critical to finding better solutions.
@@
To illustrate this, we consider the problem of removing pairs of consecutive blocks of Llama-3-8B. We perform a brute-force search over all possible $8008$ block removal configurations, where six pairs of blocks are removed. Our method identifies the optimal configuration by the 6th generation, having evaluated only 16 candidates for initialization and 8 candidates per generation. Figure 1 illustrates how the algorithm progressively approaches the optimum in terms of Hamming distance.
Consequently, EvoPress is heavily exploitation-focused: we rely on elitism, introduce minimal mutation, maintain only a single offspring, and therefore employ zero population diversity. We present ablations and a short discussion on these choices in Appendix B. EvoPress excels at optimizing smooth fitness environments, a capability we theoretically support by proving rapid convergence under an $\ell_{1}$-constraint for the class of linear functions [^1] In our context, one bit corresponds to two compression levels, while each weight of the linear function corresponds to the “saliency”. The $\ell_{1}$-constraint is now equivalent to an compression constraint..
Theorem 1.
@@ Discussion.
The proof is based on stochastic drift analysis and can be found in Appendix A. The derived bound has important practical implications. By increasing the number of offspring per generation, we can reduce the number of generations required for convergence, with the reduction scaling proportionally to $\lambda$ up to a reasonably large value. Since our approach uses a highly aggressive form of multi-step selection, the benefit is not simply a zero-sum trade-off. Evaluating many offspring in each generation incurs a significantly lower per-offspring computational cost, leading to a substantial speedup in convergence time. This makes the algorithm highly effective in smooth fitness environments, making it particularly well-suited for dynamic model compression.
@@ 4 Experiments
We now validate the efficiency of EvoPress for determining the optimal layer-wise compression across three approaches: (1) layer dropping, where the goal is to isolate the “optimal” set of blocks to drop given a target ratio, (2) non-uniform unstructured sparsity and (3) non-uniform quantization, where we are given a set of compression options per layer (sparsities or bit-widths), and the goal is to find the “optimal” configuration that matches a certain model size. We focus on LLM compression, given the major interest in the reduction of their model size and inference latency, but our method is general and can be applied to any neural network architecture and application domain.
@@
Concretely, our algorithm works as follows: Initially, for the case of quantization between available bit widths (e.g. 2.5 bit) and block dropping, we produce a number of initial configurations (around 32), evaluate them on a few data samples, and take the fittest one. For quantization with available target bitwidth and unstructured sparsity, we simply initialize using the uniform configuration. Then, we generate new offspring in each generation by making a small number of random switches in compression levels, where the number of switches is sampled from min(randint(1,3), randint(1,3)) and compression levels are exchanged in such a way that the overall compression ratio is maintained. We perform selection in multiple steps by iteratively choosing only the best configurations for survival, where each round uses progressively more tokens and has fewer survivors. To ensure elitism, we add the current parent to the candidate pool in the last stage of selection. Finally, after two or three of such stages, we take the last remaining configuration and adopt it as the population for the next round. We run for a fixed number of iterations, chosen so that performance on held-out data no longer improves. The detailed parameter setting is described in Appendix C.
To perform per-layer compression via unstructured sparsity and quantization, we adopt the data-aware compression methods SparseGPT (Frantar & Alistarh, 2023) and GPTQ (Frantar et al., 2022). For this purpose, we use Fineweb-Edu (Penedo et al., 2024) as a source of clean and diverse calibration data. Following Egiazarian et al. (2024), we fix the total number of calibration tokens to 8 million (8M). For a fair comparison, all competitive methods employ the same calibration data. The code and all configurations found by EvoPress are available at https://github.com/IST-DASLab/EvoPress.
@@
We follow the standard LLM evaluation protocol from Frantar et al. (2022). Specifically, we measure perplexity on the WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2019) datasets for language performance and accuracy on zero-shot evaluations on standard benchmarks: WinoGrande (Sakaguchi et al., 2021), PiQA (Tata & Patel, 2003), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018) via the LM Eval Harness (Gao et al., 2021).

### v2

# EvoPress: Accurate Dynamic Model Compression via Evolutionary Search

@@ Abstract
The high computational costs of large language models (LLMs) have led to a flurry of research on LLM compression, via methods such as quantization, sparsification, or structured pruning. A new frontier in this area is given by dynamic, non-uniform compression methods, which adjust the compression levels (e.g., sparsity) per-block or even per-layer in order to minimize accuracy loss, while guaranteeing a global compression threshold. Yet, current methods rely on estimating the “importance” of a given layer, implicitly assuming that layers contribute independently to the overall compression error. We begin from the motivating observation that this independence assumption does not generally hold for LLM compression: pruning a model further may even significantly recover performance. To address this, we propose EvoPress, a novel evolutionary framework for dynamic LLM compression. By formulating dynamic compression as a general optimization problem, EvoPress identifies optimal compression profiles in a highly efficient manner, and generalizes across diverse models and compression techniques. Via EvoPress, we achieve state-of-the-art performance for dynamic compression of Llama, Mistral, and Phi models, setting new benchmarks for structural pruning (block/layer dropping), unstructured sparsity, and quantization with dynamic bitwidths. Our code is available at https://github.com/IST-DASLab/EvoPress.
Machine Learning, ICML
@@ 1 Introduction
Table 1: Depth pruning is not monotone. In this example (Llama-3-8B), removing strictly more blocks (depicted in orange) can improve perplexity across sources. The left half of a block corresponds to the attention layer, the right half to the MLP.
[…]
Llama-3-8B                                                                                                                           5.54 8.80 7.72
                                                                                                                          188.01 147.25 70.46
                                                                                                                          24.39 35.53 26.24
Compression has become a standard way of reducing the deployment costs of large language models (LLMs). Current post-training techniques can be roughly categorized into quantization-based, which reduce the bit-width of weights or activations, e.g. (Frantar et al., 2022; Lin et al., 2023; Dettmers & Zettlemoyer, 2022; Tseng et al., 2024), pruning-based, which sparsify the weight matrices, e.g. (Frantar & Alistarh, 2023; Yin et al., 2024), or structured pruning / layer dropping, which drop entire model components, e.g. (Kim et al., 2024; Men et al., 2024). While improvements are still being made, existing methods are reaching diminishing returns in terms of accuracy-vs-compression (Dettmers et al., 2023; Tseng et al., 2024).
In this context, a new direction is dynamic, or non-uniform, layer-wise compression, in which different layers can be compressed to various levels, according to their “sensitivity” relative to the model output. Dynamic compression allows to maximize model accuracy while satisfying a given compression requirement, e.g. a target model size. Instance-specific solutions for this problem have already been proposed for essentially every compression type: sparsity (Yin et al., 2024), quantization (Frantar & Alistarh, 2022), or layer dropping (Kim et al., 2024; Men et al., 2024). Broadly, these approaches work by assigning an error/sensitivity score to each layer and compression level, which measures the impact of its compression on output loss increase. Then, one calculates a compression assignment which minimizes the sum of error scores, while still satisfying the global compression constraint. Thus, such approaches inherently assume error monotonicity: i.e., that a lower sum of error scores taken over layers implies a lower compression error for the entire model.
Our work starts from the observation that error monotonicity does not hold generally for LLM compression: specifically, there are instances where compressed models with lower sums of per-layer errors can perform worse than models with higher error. We illustrate this fact in Table 1, which shows an instance of a layer dropping configuration where pruning more blocks leads to significantly better perplexity than an instance which prunes strictly fewer blocks.
@@
Our approach builds on the observation that the partial effectiveness of prior work, which assumes error linearity, suggests that the fitness landscape induced by compression allocation has properties similar to those of a linear function. Such functions are particularly well-suited to hill climbing, meaning they can be optimized through highly local exploration, akin to gradient descent in a continuous search space. EvoPress is specifically designed as an efficient hill climber, though it is capable of optimizing a broad class of fitness environments. Notably, our algorithm guarantees convergence: specifically, any linear fitness function defined on the $n$-dimensional hypercube will be maximized in expected $O(k(n-k)/\lambda)$ generations under the constraint $\|x\|_{1}=k$, where $\lambda$ is the number of offspring. The proof is quite non-trivial, as it needs to adapt stochastic drift analysis techniques to the case where multiple offspring are examined in each sub-step. In Figure 1, we illustrate the algorithm’s fast convergence and high efficiency on a practical example with correlated block dropping on Llama-3-8B, where we determined the optimum via (expensive) exhaustive search: EvoPress is able to reach the optimum in only $6$ generations, using a total of only $56$ model evaluations. A key advantage of our approach is that it is agnostic of the model architecture and compression type. We illustrate this via experiments, which are the first to span all three compression methods, across different LLM families.
Results show that EvoPress significantly improves upon all prior work on depth pruning in terms of accuracy-vs-compression, especially at medium levels, and also outperforms the prior best methods - OWL and dynamic programming, respectively - for non-uniform pruning and quantization. Moreover, it can do so efficiently: the full version of EvoPress, applied at high compression granularity, will converge in a few hours on a single RTX 3090 GPU, and we also present a lightweight version which utilizes fewer samples and converges in $\sim$ 1 hour in the same setting, on an 8B-parameter model.
@@ 2 Related Work
To our knowledge, we are the first to present a unified approach which covers all types of post-training LLM compression (i.e., layer dropping / depth pruning and non-uniform pruning / quantization) - so far, these problems have generally been approached independently.
Depth Pruning. Recently, there has been a lot of interest in compression by removing entire transformer blocks, both for efficiency and to gain insights about the language model itself. Most methods are based on scoring the importance of each block, and then maximizing the importance of the resulting model by removing the blocks of lowest importance. Therefore, they assume error linearity, meaning that each block contributes independently to the total compression error. Weight Subcloning (Samragh et al., 2023) proposed a multi-step process to find good initializations for an untrained smaller model given an already trained larger one, where the importance of each block is scored based on the ratio of $\ell_{2}$ norms between the output embeddings of the block with and without the residual connection. Shortened Llama (Kim et al., 2024) proposes scoring each block by measuring the perplexity after removing the respective block from the full model. ShortGPT (Men et al., 2024) uses the cosine similarity between the input and output embeddings of each block to assess its importance. By contrast, Gromov et al. (2024) restrict to removing consecutive blocks and score each configuration by cosine similarity.
@@
The recent OWL method (Yin et al., 2024) focuses on non-uniform pruning of LLMs, and provides consistent improvements over uniform profiles via a layer scoring system which analyzes the activation outlier structure. Experimentally, we find that OWL is effective especially for Llama-family models (Touvron et al., 2023) and at moderate sparsities, but observe significant gaps in favor of EvoPress across all models and compression levels.
@@
Random search is also popular in the context of structural pruning and Neural Architecture Search (NAS) (Chen et al., 2020; Dong et al., 2021; Wang et al., 2020; Xu et al., 2021; Yin et al., 2021; Molchanov et al., 2022; Kurtić et al., 2024). However, such methods rely on re-training and have notoriously high costs, which limits their applicability to post-training compression. Thanks to its low sample complexity, we believe that EvoPress could be extensible to lightweight NAS, and plan to investigate this in future work.
@@ 3 Method
where $f(\hat{M})$ quantifies the performance of the compressed model $\hat{M}$ and $g(\hat{M})$ represents the compression constraint. For simplicity, we will define $g$ as the model’s total size (in terms of parameters); however, the method can be adapted to other constraints, such as inference speed.
@@
Given the specified database, any compressed model is characterized by its compression level per unit (e.g. per layer). With $n$ units, each available in $m$ compression levels, our objective is to find
@@
where we are searching over the set of $n$-tuples over $[m]$. Assessing the performance of a model in practice typically involves benchmark tasks, which have limited scope and require lengthy evaluation. We address these challenges by using the base model as the gold standard and focusing solely on the relative degradation of our compressed models. To quantify this degradation, we measure the Kullback-Leibler (KL) divergence between the two models, as it has proven particularly robust with limited data. Empirically, we observed that already around 64K tokens of calibration data (corresponding to $8$ full sample sequences for Llama-3-8B) are sufficient to reliably determine the quality of the lightweight model. To avoid confusion, we will refrain from inverting the fitness function and from now on consider the minimization problem
$\hat{M}^{*}=\operatorname*{arg\,min}_{v\in[m]^{n}}D_{KL}(P_{M}\parallel Q_{\hat{M}_{v}})\;\,\text{subject to}\;\,g(\hat{M_{v}})\leq C,$
@@
Initialization: $\texttt{candidates}\leftarrow[]$ ;
for $i\leftarrow 1$ to initialCandidates do
       $\texttt{candidate}\leftarrow\texttt{sampleUniformly}()$;
@@
$x^{(1)}\leftarrow\texttt{selectTopKFittest}(\texttt{candidates},\newline \texttt{initialTokens},K=1)$;
@@
             $y_{i}\leftarrow\texttt{LevelSwitchMutation}(y_{i})$;
             $\texttt{offspring}.\texttt{append}(y_{i})$;
       Selection: for $\texttt{step}\leftarrow 1$ to selectSteps do
@@
       $x^{(t+1)}\leftarrow\texttt{offspring}[0]$;
Algorithm 1 EvoPress: A $(1+\lambda)$-Evolutionary Algorithm with Level-Switch Mutation and Multi-Step Selection for Maximizing $f:[m]^{n}\rightarrow\mathbb{R}$.
@@
For initialization, we apply the target level directly if it matches an available setting (e.g., all layers at 70% sparsity for an average of 70% sparsity). If the target falls between two compression levels (e.g., for block dropping), we initialize by randomly sampling candidates with some units compressed to the next lower level, and others to the next higher level, selecting the fittest among them. A high level overview of this procedure can be found in Algorithm 1.
@@
Crossover, for instance, is only effective if population diversity is preserved, for example measured by the sum of pairwise Hamming distances between individuals (Jansen & Wegener, 2002; Opris et al., 2024). While this promotes a more thorough exploration of the search space, it requires allocating resources to less promising regions, which slows down the progress toward optimal solutions. Similarly, non-elitist algorithms, despite their ability to escape local optima (Dang et al., 2021; Jorritsma et al., 2023; Lengler et al., 2024), also incur costs by frequently discarding potentially useful individuals. Consequently, these approaches should be reserved for situations where the fitness landscape is rugged, and escaping local optima is critical to finding better solutions.
@@
To illustrate this, we consider the problem of removing pairs of consecutive blocks of Llama-3-8B. We perform a brute-force search over all possible $8008$ block removal configurations, where six pairs of blocks are removed. Our method identifies the optimal configuration by the 6th generation, having evaluated only 16 candidates for initialization and 8 candidates per generation. Figure 1 illustrates how the algorithm approaches the optimum in Hamming distance.
Consequently, EvoPress is heavily exploitation-focused: we rely on elitism, introduce minimal mutation, maintain only a single offspring, and therefore employ zero population diversity. We present ablations and a short discussion on these choices in Appendix B. EvoPress excels at optimizing smooth fitness environments, a capability we theoretically support by proving rapid convergence under an $\ell_{1}$-constraint for the class of linear functions. (Here, one bit corresponds to two compression levels, while each weight of the linear function corresponds to the “saliency”. The $\ell_{1}$-constraint is now equivalent to a compression constraint.)
Theorem 3.1.
@@ Discussion.
The proof is based on stochastic drift analysis and can be found in Appendix A. Notably, by increasing the number of offspring per generation, we can reduce the number of generations required for convergence, with the reduction scaling proportionally to $\lambda$ up to a reasonably large value. Since our approach uses a highly aggressive form of multi-step selection, the benefit is not simply a zero-sum trade-off. Evaluating many offspring in each generation incurs a significantly lower per-offspring computational cost, leading to a substantial speedup in convergence time. This makes the algorithm highly efficient in relatively smooth fitness environments.
@@ 4 Experiments
We now validate the effectiveness of EvoPress for determining the optimal layer-wise compression across three approaches: (1) layer dropping, where the goal is to isolate the “optimal” set of blocks to drop given a target ratio, (2) non-uniform unstructured sparsity and (3) non-uniform quantization, where we are given a set of compression options per layer (sparsities or bit-widths), and the goal is to find the “optimal” configuration that matches a certain model size. We focus on LLM compression, given the major interest in the reduction of their model size and inference latency, but our method is general and can be applied to any neural network architecture and application domain.
@@
Concretely, our algorithm works as follows: Initially, for the case of quantization between available bit widths (e.g. 2.5 bit) and block dropping, we produce a number of initial configurations (around 32), evaluate them on a few data samples, and take the fittest one. For quantization with available target bitwidth and unstructured sparsity, we simply initialize using the uniform configuration. Then, we generate new offspring in each generation by making a small number of random switches in compression levels, where the number of switches is sampled from min(randint(1,3), randint(1,3)) and compression levels are exchanged in such a way that the overall compression ratio is maintained. We perform selection in multiple steps by iteratively choosing only the best configurations for survival, where each round uses progressively more tokens and has fewer survivors. To ensure elitism, we add the current parent to the candidate pool in the last stage of selection. Finally, after two or three of such stages, we take the last remaining configuration and adopt it as the population for the next round. We selected the number of generations, offspring, and tokens based on the search space size but found the search to be highly robust. Even drastic changes in hyperparameter settings yield similar results (see, e.g., Figure 3). The detailed parameter setting is described in Appendix C.1. For our main results we used a fixed number of generations for optimization, which was chosen conservatively to better understand convergence behavior. In practice, stopping the search early can significantly reduce runtime requirements. We discuss a simple stopping criteria and provide runtime comparisons in Appendix G.
To perform per-layer compression via unstructured sparsity and quantization, we adopt the data-aware compression methods SparseGPT (Frantar & Alistarh, 2023) and GPTQ (Frantar et al., 2022). For this purpose, we use Fineweb-Edu (Penedo et al., 2024) as a source of clean and diverse calibration data. Following Egiazarian et al. (2024), we fix the total number of calibration tokens to 8 million (8M). For a fair comparison, all competitive methods employ the same calibration data.
@@
We follow a standard evaluation protocol (Frantar et al., 2022), measuring perplexity on the WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2019) datasets for language performance and accuracy on zero-shot evaluations on standard benchmarks: WinoGrande (Sakaguchi et al., 2021), PiQA (Tata & Patel, 2003), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018) via the LM Eval Harness (Gao et al., 2021).